On sensing danger two birds Va=i+4j +3k and Vb=4i+2j-4k relative to a stationary observer. How far apart are the two bird after 5secs
1
Expert's answer
2021-08-25T12:25:54-0400

"\vec V=3\vec i-2\vec j-7\vec k,"

"V=\sqrt{3^2+(-2)^2+(-7)^2}=7.87~\frac ms,"

"s=Vt=39.4~m."